# DGSNA: prompt-based Dynamic Generative Scene-based Noise Addition method

### v2

@@ Abstract
noise addition method , scene-based noise , scene information[label1]organization=School of Computer Science and Technology, Guangdong University of Technology,addressline=No.100 Waihuan West Road, city=Guangzhou, postcode=510006, state=Guangdong, country=China
@@ 4.5 Experimental parameters
Additionally, within the TTA system employed in this experiment, the chosen model was AudioLDM-l-full [42]. The number of diffusion steps (ddim
_steps) for the Denoising Diffusion Implicit Model (DDIM) [61] was set to 200, controlling the iterations of the denoising process. A guidance scale (guidance_scale) of 2.5 was employed to steer the model’s output based on provided guidance, enhancing alignment with target audio characteristics. The generated audio duration (duration) was set to 5.0 seconds.
@@ 5 Experimental results
Model | Train Set ANR | Stream Decode | Test WER() | |||||||
Aishell-1 ANR | WenetSpeech | |||||||||
0 | 10 | 20 | 30 | 40 | Dev | Test_Net | Test_Meeting | |||
Baseline | 0% | True | 5.33 | 9.90 | 14.30 | 19.04 | 23.72 | 29.90 | 48.94 | 53.99 |
False | 4.93 | 9.47 | 13.86 | 18.51 | 23.23 | 29.16 | 48.29 | 53.60 | ||
DGSNA_10 | 10% | True | ||||||||
False | ||||||||||
DGSNA_20 | 20% | True | ||||||||
False | ||||||||||
DGSNA_30 | 30% | True | ||||||||
False | ||||||||||
DGSNA_40 | 40% | True | ||||||||
False |
@@ 5.2 Comparison of different ANR
Table 2 presents the ASR experimental results. Introducing DGSNA during training significantly improves performance in both known noisy scenarios (Aishell-1 with ANR 0) and, to a lesser extent, unknown scenarios (WenetSpeech), achieving up to 11.21% relative improvement. This improvement comes at the cost of slight performance degradation in clean scenarios (Aishell-1 with ANR = 0). The DGSNA application rate (ANR) significantly influences results. We hypothesize that a low ANR (10%) results in insufficient data augmentation, while a high ANR (40%) leads to overfitting. An ANR of 20% proved optimal.
@@ 5.3.1 BET prompt
To ensure compatibility across various model architectures, it is essential to tailor the BET prompt according to the operational mode of the target generative chat model. There are two common classifications of generative chat models based on parameter-passing methods: (a) Single Parameter Input Models: In this configuration, models like BlueLM-Chat concatenate both the current chat input and historical chat information into a single consolidated parameter, often referred to as the prompt. (b) Dual Parameter Input Models: Contrarily, models such as ChatGLM3 are designed to handle the current chat input and historical chat information as two separate parameters, denoted as the prompt and history. As demonstrated in Figure 7, we have developed tailored prompts for two distinct models, BlueLM-Chat and ChatGLM3, corresponding to their respective single and dual parameter input frameworks.
@@ 5.4.1 Correct Example
This subsection offers an in-depth analysis of the workflow for generating scene-based speech, specifically for the sample Qwen-7B-Chat_pedestrian_street
_1894.wav, as shown in Figure 9. This process highlights the comprehensive steps taken to ensure the production of high-quality, contextually appropriate scene-based speech using the DGSNA.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by the National Science Foundation of China Under Grant U21A20478, and in part by the 2022 Industrial Technology Basic Public Service Platform Project of China (No.2022-228-219), and in part by Key Laboratory of MIIT for Intelligent Products Testing and Reliability 2023 Key Laboratory Open Project Fund (No. CEPREI2023-01).

### v5

@@ Abstract
noise addition method , scene-based noise , scene-based information[label1]organization=School of Computer Science and Technology, Guangdong University of Technology,addressline=No.100 Waihuan West Road, city=Guangzhou, postcode=510006, state=Guangdong, country=China
@@ 4.5 Experimental parameters
Additionally, within the TTA system employed in this experiment, the chosen model was AudioLDM-l-full [42]. The number of diffusion steps (ddim_
steps) for the Denoising Diffusion Implicit Model (DDIM) [61] was set to 200, controlling the iterations of the denoising process. A guidance scale (guidance_scale) of 2.5 was employed to steer the model’s output based on provided guidance, enhancing alignment with target audio characteristics. The generated audio duration (duration) was set to 5.0 seconds.
@@ 5.2 Comparison of different ANR
Model | Train Set ANR | Stream Decode | Test WER() | ||||
Aishell-1 ANR | |||||||
0 | 10 | 20 | 30 | 40 | |||
Baseline | 0% | True | 5.33 | 9.90 | 14.30 | 19.04 | 23.72 |
False | 4.93 | 9.47 | 13.86 | 18.51 | 23.23 | ||
DGSNA_10 | 10% | True | |||||
False | |||||||
DGSNA_20 | 20% | True | |||||
False | |||||||
DGSNA_30 | 30% | True | |||||
False | |||||||
DGSNA_40 | 40% | True | |||||
False |
Model | Train Set ANR | Stream Decode | Test WER() | |||
Aishell-1 ANR | WenetSpeech | |||||
0 | Dev | Test_Net | Test_Meeting | |||
Baseline | 0% | True | 5.33 | 29.90 | 48.94 | 53.99 |
False | 4.93 | 29.16 | 48.29 | 53.60 | ||
DGSNA_10 | 10% | True | ||||
False | ||||||
DGSNA_20 | 20% | True | ||||
False | ||||||
DGSNA_30 | 30% | True | ||||
False | ||||||
DGSNA_40 | 40% | True | ||||
False |
Table 2 and Table 3 presents the ASR experimental results. Introducing DGSNA during training significantly improves performance in both known noisy scenarios (Aishell-1 with ANR 0) and, to a lesser extent, unknown scenarios (WenetSpeech), achieving up to 11.21% relative improvement. This improvement comes at the cost of slight performance degradation in clean scenarios (Aishell-1 with ANR = 0). The DGSNA application rate (ANR) significantly influences results. We hypothesize that a low ANR (10%) results in insufficient data augmentation, while a high ANR (40%) leads to overfitting. An ANR of 20% proved optimal.
@@ 5.3.1 BET prompt
To ensure compatibility across diverse model architectures, tailoring the BET prompt to the operational mode of the target generative chat model is crucial. Generative chat models are commonly classified into two types based on their parameter-passing methods: (a) Single Parameter Input Models: In this configuration, models such as BlueLM-Chat concatenate the current chat input and historical chat information into a single consolidated parameter, typically termed the prompt. (b) Dual Parameter Input Models: Conversely, models like ChatGLM3 are designed to process the current chat input and historical chat information as two separate parameters, referred to as the prompt and history. As illustrated in Figure 7, we developed tailored prompts for two distinct models, BlueLM-Chat and ChatGLM3, corresponding to their respective single and dual parameter input frameworks.
@@ 5.4.1 Correct Example
This subsection provides a detailed analysis of the workflow for generating scene-based speech, using the sample depicted in Figure 9 as an illustration. This process underscores the comprehensive steps undertaken to ensure the production of high-quality, contextually appropriate scene-based speech utilizing DGSNA.
@@ 6 Conclusion
Acknowledgments
This work is supported in part by National Research and Development Key Project No. 2023YFF0905603, and in part by the 2022 Industrial Technology Basic Public Service Platform Project of China (No.2022-228-219), and in part by Key Laboratory of MIIT for Intelligent Products Testing and Reliability 2023 Key Laboratory Open Project Fund (No. CEPREI2023-01).